Accelerator Update for January 7, 2008 - January 9, 2008
Monday January 7
The midnight shift began with Operations monitoring the Tevatron (TeV) store 5830 (52.45E30) and the Recycler stash (391.9E10), with the Antiproton Source stacking (59.6E10), with the Linac and Booster operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the NuMI and SciBooNE / MiniBooNE (BNB) experiments taking data.
Operations terminated store 5830 at 8:45 AM.
D0 experimenters accessed their collision hall at 9:30 AM.
A MI expert began a one-hour coalescing study at 9:32 AM.
A MI expert began a collimator study at 11:36 AM.
Vacuum technicians came out of the TeV at 12:20 PM.
D0 came out of access at 12:24 PM.
The MI slipstacking study ended at 12:48 PM.
Pbar resumed stacking at 1:04 PM.
Booster experts began an orbit monitor study at 1:13 PM.
Pbar experts began a reverse proton study at 1:23 PM.
At 2:12 PM, a MI RF station (MIRF2) tripped off for the first of many times in 13 minutes. It reset after each trip. Operators contacted an RF expert.
At 3:09 PM, Operations reported that the TeV sector A49 vacuum pump tripped off. An expert investigated and reported that its wall breaker had tripped off and would need to be replaced.
Pbar experts completed their reverse proton study at 3:15 PM.
At 3:20 PM, MIRF2 tripped off and caused the NuMI beam permit to trip off as well. Everything reset.
Operations put the TeV into shot setup at 3:41 PM.
Operations established store 5832 at 6:20 PM with an initial luminosity of 249.54E30.
At 6:38 PM, a Linac Klystron RF station (KRF2) tripped off. It reset.
At 6:42 PM, MIRF8 began tripping off with every pulse. Operators put in the cavity short and toggled the NuMI beam permit input, but this didn’t help. They contacted and expert who helped them bring the station back on line.
Pbar and NuMI resumed taking beam at 7:32 PM.
Operations reported at midnight that the Recycler had stashed antiprotons three times during the evening shift.
Tuesday January 8
The midnight shift began with Operations monitoring store 5832 (115.66E30) and the stash (184.4E10), and with NuMI and BNB taking beam.
At 12:37 AM the 400MeV area beam permit tripped off due to bad beam positions. It reset.
At 1:26 AM, LRF4 tripped off. It reset.
At 1:39 AM, KRF2 tripped off. It reset.
At 2:10 AM, KRF1 tripped off. It reset.
At 2:18 AM, a TeV sector A0 vacuum pump tripped off. Operators will reset this between stores.
At 5:07 AM, BRF14 tripped off. It reset.
Cryo system experts reported at 7:43 AM that the wet engine failure was due to and electrical fault.
Operations reported at 7:54 AM that the Recycler had stashed four times during the midnight shift.
Operations prepared D0 for an access at 9:28 AM.
The Cryo Coordinator reported at 9:42 AM that the regeneration unit for the A1 wet engine needed to be replaced.
A separator expert reported at 9:59 AM that he would be working on the TeV C17 separator.
An instrumentation expert began working on the D0 BLM chassis at 10:18 AM.
The Cryo Coordinator reported at 11:16 AM that cryo techs had completed their work on the A1 wet engine.
At 11:16 AM, Operations reported that a Linac quadrupole power supply (L:QPS511) was bouncing out of tolerance. A Linac expert investigated.
The Neutron Therapy Facility (NTF) reported that they had completed treatment on their last patient for the day at 12:52 PM.
Operations put the TeV through a dry squeeze at 1:04 PM.
At 1:10 PM, Operations reported a problem with a Pbar kicker (PBKICK) and two MI sweepers. A reset didn’t work. Operations contacted a Pbar expert who had everything back on five minutes later.
Operations reported at 1:50 PM that safety system experts were beginning work on the TeV sector E4 safety system.
Operations put the TeV through a dry squeeze at 2:47 PM.
Operations put the TeV into shot setup at 3:16 PM.
Operations reported at 4:06 PM that the Recycler had stashed four times during the day shift.
Operations prepared the Pelletron for access at 4:47 PM.
Operations established store 5835 at 5:08 PM with an initial luminosity of 259.92E30.
Operators and Recycler experts began a cooling study at 6:02 PM.
At 7:38 PM, Operations reported that a Linac Klystron LCW water skid had been in alarm on low pressure all shift. Operators contacted experts who said that this was the result of an MTA Klystron that had been valved in and it was okay.
Wednesday January 9
The midnight shift began with Operations monitoring store 5835 (110.67E30) and the stash (13.9E10), with Pbar stacking (171.7E10), and with NuMI and BNB taking beam.
At 5:25 AM, MIRF8 began tripping off. Operations turned off MI.
Operations put Pbar into standby at 5:31 AM.
Operations began, at 5:56 AM, to prepare NuMI, Booster, and MI for a scheduled access.
The Plan for Wednesday
The plan for today is to conduct specific jobs in the Booster and the Main Injector. The Booster work should be completed by noon and the MI work by early evening. The Run Coordinator would like to see a shot in the TeV by midnight. Pbar set a new stacking record with a rate of 24.67E10 for one hour.
Comments and Suggestions
|last modified 01/09/2008 email Fermilab|
|Security, Privacy, Legal|